OpenStudy (anonymous):

What is the integral of e^9x

OpenStudy (shayaan_mustafa):

I[e^(9x)dx] Replace 9x with u to simplify the integrand. u=9x Find an equation relating the value of dx to du. du=9 dx Since du=9 dx, then dx=(du)/((9)). dx=(du)/(9) Replace the value found for dx so that the integrand is in terms of u. I[,,e^(u)*(1)/(9),u] Simplify the integrand. I[,,(e^(u))/(9),u] The integral of (e^(u))/(9) is (e^(u))/(9). (e^(u))/(9)+C Replace u=9x in the solved integral to find the solution in terms of x. (e^(9x))/(9)+C NOTE: I=Integral

OpenStudy (shayaan_mustafa):

after comma. it is du (,u)

OpenStudy (anonymous):

I'm kind of confused at the part where you put the integrand in terms of x. Why is there a second u at the end (after the comma)?

OpenStudy (shayaan_mustafa):

where? line no. please.